# Linking Process to Outcome: Conditional Reward Modeling for LLM Reasoning

### v1

@@ Appendix A Implementation Details
The code and checkpoints will be released upon the acceptance of this paper.
@@ A.4 RL Optimization
where $\pi_{\theta_{\text{old}}}$ is old policy and $\pi_{\theta_{\text{ref}}}$ is the reference model. $\epsilon$ is the PPO clipping threshold and $\beta$ controls the strength of the KL penalty; both are hyperparameters.
@@ Appendix E Cross-sample Comparison Capability
The core modeling principle of PQM (Li & Li, 2025) focuses on the relative ordering of steps within a sample, rather than modeling absolute values. For instance, applying the same shift operation to the $Q$-values of all steps within a sample still yields the same optimal ordering. As a result, the reward magnitudes across different samples do not share a consistent meaning. As stated in Corollary D.1 of the PQM paper, PQM allows comparison only when two trajectories share the same correct prefix. The limitation is validated by its suboptimal performance in our Best-of-N and beam search experiments (Section 4.2, Section 4.3), further demonstrating PQM’s deficiency in cross-sample comparability.
@@
A core deficiency of IPRM (Yuan et al., 2025) is that its training relies solely on the final outcome, which leads to an ambiguous credit assignment, as the model only ensures that the aggregate of all step-level rewards aligns with the final result, without enforcing more rigorous constraints on the process. Consequently, it lacks explicit modeling of the fine-grained causal relationships between reasoning steps. The process reward for any given step is not necessarily coherent with the preceding trajectory, resulting in inconsistent process rewards that cannot be reliably compared across different samples.
@@ Appendix F Case studies on Process Rewards
height=1.8ex]1 1 The ORM fails to provide meaningful process rewards. Its rewards are volatile and poorly aligned with the correctness of intermediate steps. 2 The PQM, while capturing local ranking dependencies between reasoning steps, receives only indirect supervision from the final outcome. Consequently, it may incorrectly assign high rewards after a fatal error, failing to provide a precise assessment of the final result. 3 In contrast, our CRM overcomes these issues and precisely identifies which reasoning step enters an incorrect state. It maintains high rewards for correct steps, exhibits an immediate drop when the first error occurs, and subsequently maintains a low reward for all following incorrect steps.

### v2

@@ 5 Conclusion
6 Acknowledgments
The research received partial support from National Natural Science Foundation of China (Grant No. 62406193). The authors also gratefully acknowledge further assistance provided by Shanghai Frontiers Science Center of Human-centered Artificial Intelligence, MoE Key Lab of Intelligent Perception and Human-Machine Collaboration, and HPC Platform of ShanghaiTech University.
@@ A.4 RL Optimization
where $\pi_{\theta_{\text{old}}}$ is the old policy and $\pi_{\theta_{\text{ref}}}$ is the reference model. $\epsilon$ is the PPO clipping threshold and $\beta$ controls the strength of the KL penalty; both are hyperparameters.
@@
CRM+VR setting.
When used CRM alone, it provides step-wise rewards ($r_{t}^{C}$) for all steps, including the final reasoning step $T$. When combined with VR, only the final step additionally receives a verifiable reward ($r_{T}^{V}$), while earlier steps still rely solely on CRM. Formally, the reward at each step $t$ is defined as:
$r_{t}=\begin{cases}r_{t}^{C},&\text{if }t<T\\ r_{T}^{C}+r_{T}^{V},&\text{if }t=T\end{cases}$ (17)
@@ Appendix E Cross-sample Comparison Capability
The core modeling principle of PQM (Li and Li, 2025) focuses on the relative ordering of steps within a sample, rather than modeling absolute values. For instance, applying the same shift operation to the $Q$-values of all steps within a sample still yields the same optimal ordering. As a result, the reward magnitudes across different samples do not share a consistent meaning. As stated in Corollary D.1 of the PQM paper, PQM allows comparison only when two trajectories share the same correct prefix. The limitation is validated by its suboptimal performance in our Best-of-N and beam search experiments (Section 4.2, Section 4.3), further demonstrating PQM’s deficiency in cross-sample comparability.
@@
A core deficiency of IPRM (Yuan et al., 2025) is that its training relies solely on the final outcome, which leads to ambiguous credit assignment, as the model only ensures that the aggregate of all step-level rewards aligns with the final result, without enforcing more rigorous constraints on the process. Consequently, it lacks explicit modeling of the fine-grained causal relationships between reasoning steps. The process reward for any given step is not necessarily coherent with the preceding trajectory, resulting in inconsistent process rewards that cannot be reliably compared across different samples.
@@ Appendix F Case Studies on Process Rewards
1
The ORM fails to provide meaningful process rewards. Its rewards are volatile and poorly aligned with the correctness of intermediate steps. 2 The PQM, while capturing local ranking dependencies between reasoning steps, receives only indirect supervision from the final outcome. Consequently, it may incorrectly assign high rewards after a fatal error, failing to provide a precise assessment of the final result. 3 In contrast, our CRM overcomes these issues and precisely identifies which reasoning step enters an incorrect state. It maintains high rewards for correct steps, exhibits an immediate drop when the first error occurs, and subsequently maintains a low reward for all following incorrect steps.
@@ Appendix G Ablation Study on Process Reward Formulation of CRM
Appendix H Comparison with Reasoning-enhanced PRMs
Table 11: Best-of-N accuracy comparison with reasoning-enhanced PRMs on MATH500.
Method Best-of-8 Best-of-16
R-PRM 49.8 51.2
GenPRM 50.8 55.2
CRM (ours) 53.0 56.4
Recent work (Chen et al., 2025; Guo et al., 2025b; She et al., 2025; Zhao et al., 2025) integrates explicit reasoning capabilities into reward modeling. To compare our CRM with this line of work, we evaluate it against the representative reasoning‑enhanced reward model R‑PRM (She et al., 2025) and GenPRM (Zhao et al., 2025).
We evaluated these reward models using their official released checkpoints, the same settings and inference pipelines to score identical responses from our generator on the MATH500 dataset under Best-of-N sampling. The results are shown in Table 11. Compared to both baselines, CRM consistently achieves the highest performance. Specifically, under the Best-of-8 setting, CRM achieves 53.0, surpassing R-PRM (49.8) and GenPRM (50.8). Similarly, under the Best-of-16 setting, CRM reaches 56.4, outperforming both R-PRM (51.2) and GenPRM (55.2). These results highlight CRM’s stronger capability in identifying high-quality responses compared to recent reasoning-enhanced and generative approaches.
Appendix I Best-of-N sampling performance characteristics
In this section, we provide a more in-depth analysis of the performance characteristics of Best-of-N sampling. In BoN, an LLM generator first samples $N$ candidates for each query. Subsequently, a reward model selects the response with the highest predicted reward. Therefore, the LLM generator plays a role in shaping BoN performance. We observe that different LLM generators lead to distinct behaviors: BoN performance may continue to improve as $N$ increases, or it may converge at a small $N$ and show no further gains. As shown in Table 12, with a strong generator (Llama-3.1-8B-Instruct), all methods converge at small $N$ and do not improve further, whereas with a weaker generator (MetaMath-Mistral-7B), all methods exhibit increasing BoN performance as $N$ increases. This phenomenon can be explained as follows. (i) Weaker LLM generators have a lower probability of producing correct responses. At small $N$, they may generate no correct candidates, leaving the reward model with nothing correct to select. As $N$ increases, the sampled candidates are more likely to include correct answers, so the BoN performance starts to improve. (ii) Stronger LLM generators are able to produce correct responses even with small $N$. In this case, the BoN performance is not limited, leading to faster convergence.
These points are further supported by the Pass@N metric, which measures the probability that the LLM generator produces at least one correct response among the $N$ responses, serving as an upper bound on the achievable performance of BoN sampling with any reward model. Table 13 reports the Pass@N of the two LLM generators. The weaker generator MetaMath-Mistral-7B shows a substantial gap between Pass@8 and Pass@128, with a relatively low Pass@8. This indicates that when $N$ is small, the low upper bound limits BoN performance. The stronger LLM generator LLaMA-3.1-8B-Instruct achieves high Pass@N even for small $N$, leading to early convergence.
Table 12: Best-of-N sampling performance across methods using different LLM generators.
LLM Generator Method MATH500
@8 @16 @32 @64 @128
MetaMath-Mistral-7B Pass@N 53.0 62.0 69.6 75.8 80.4
PRM 34.0 35.8 35.2 37.4 38.0
PQM 36.4 38.4 37.8 38.6 40.4
CRM (ours) 35.6 39.0 39.2 40.2 41.6
LLaMA-3.1-8B-Instruct Pass@N 73.8 79.2 84.6 87.4 90.2
PRM 54.2 55.2 55.2 54.2 54.6
PQM 53.2 54.4 54.8 54.8 55.8
CRM (ours) 53.0 56.4 56.6 55.8 56.6
Table 13: Pass@N of different LLM generators on MATH500 and GSM-Plus datasets.
LLM Generator MATH500 (Pass@N) GSM-Plus (Pass@N)
[…]
MetaMath-Mistral-7B 27.4 53.0 62.0 69.6 75.8 80.4 11.1 73.2 78.1 80.5 83.0 84.3
LLaMA-3.1-8B-Instruct 16.4 73.8 79.2 84.6 87.4 90.2 6.3 79.2 81.2 82.2 83.3 85.5
Appendix J Data Efficiency Analysis of CRM.
To further substantiate the data efficiency of CRM, we added PRM for comparison. We refer to the data from the PRM training set, excluding the final step, as process data. We conducted experiments using the same data proportion as in our ablation study, except for 0% (since PRM requires process data). As shown in Table  14, CRM reaches near-optimal performance with only 50% of the training data, whereas PRM still exhibits a noticeable performance gap under the same data budget.
The efficiency gain stems from how CRM leverages the supervision at step $t=z$. While PRM treats each step independently and learns a step-wise correctness classifier, CRM instead models reasoning as a temporal progression toward the final answer. By framing the entire reasoning trajectory as a sequence converging to correctness, as derived in Eqs. 11, 12, and 13, CRM allows the supervision signal at step $t=z$ to propagate backward to all earlier steps $t^{\prime}\leq z$. This provides richer and more coherent training signals and enables CRM to make more effective use of available data, resulting in significantly improved data efficiency.
Table 14: Best-of-N accuracy on Math500 using varying proportions of data for CRM’s $\mathcal{L}_{z}$ and PRM.
Proportion of data used for L_zL\_z CRM Process data used for PRM PRM
@8 @16 @32 @64 @128 @8 @16 @32 @64 @128
0% 47.0 44.2 41.6 39.2 38.2 0% - - - - -
10% 52.4 51.2 50.6 49.6 47.6 10% 48.0 49.4 48.6 48.0 47.4
25% 54.0 52.4 53.6 53.2 52.0 25% 53.0 54.0 54.6 52.6 50.8
50% 54.4 53.6 57.2 55.8 55.0 50% 52.4 54.2 55.2 53.4 52.0
100% 53.0 56.4 56.6 55.8 56.6 100% 54.2 55.2 55.2 54.2 54.6